# Geometric Freeze-Tag Problem

### v2

@@ 4.1. Our strategy for FTP in $(\mathbb{R}^{3},l_{1})$
Given a set of $n$ asleep robots, our wake-up strategy is as follows: If $n\leq 127$, the awake robot at the origin can activate all robots within $13r$ time units, as shown in Lemma 4.
If $n>127$, by Lemma 4, we first wake up 7 robots in $3r$ time units. Then, we divide the cross-polytope into six smaller overlapping cross-polytopes, each with a diameter of $\frac{2d}{3}$, as described in Lemma 4.
With eight robots awake, we assign six of them to one of the smaller cross-polytopes. Each robot reaches the center of its assigned cross-polytope in at most $\frac{2d}{3}$ time units. Then, each robot wakes up the remaining robots in its section, taking at most $f\left(\frac{2d}{3}\right)$ time units, where $f(d)$ is the time needed to wake all robots in a region of diameter $d$. If a robot belongs to multiple cross-polytopes, it can be assigned to any one of them.
If a cross-polytope has fewer than $128$ asleep robots, they can be woken up in $\frac{13d}{4}$ time units as per Lemma 4. Thus, the recursive formula for computing $f(d)$ is:
@@
Solving this recursively yields:
$f(d)\leq\frac{13d}{6}\times\left(1+\frac{2}{3}+\frac{4}{9}+\ldots\right)=\frac{13d}{2}$
For the unit $l_{1}$-ball where $d=2$, the upper bound for the total time is $13$ units.
@@ 5.1. Mapping
Consider the upper hemisphere of a unit $l_{1}$-ball with $T=(0,0,1)$. Any point $p=(x,y,z)$ on the surface of the hemisphere can be uniquely represented by the pair $(\delta,\theta)$, where $\delta=\arccos(z)$ is the geodesic distance from $T$ to $p$ (the shortest path along the hemisphere), and $\theta$ is the angle of $p$’s projection onto the $xy$-plane relative to the positive $x$-axis. This is illustrated in Figure 8.
@@
$x$$y$$z$$\theta$$\delta$$T$$p$$\delta$$p^{\prime}$$r=\frac{\pi}{2}$$\tiny{\theta}$$x$$y$
@@
Now, let’s show that $\|p_{1}-p_{2}\|_{2}\leq\|p^{\prime}_{1}-p^{\prime}_{2}\|_{2}$. Consequently we need to show,
@@ 5.2. Our strategy, when asleep robots are on the boundary on the unit $l_{2}$-ball
For an instance of FTP in $(\mathbb{R}^{3},l_{2})$, where the asleep robots are on the boundary and the initially active robot is at the center of the ball, we propose the following strategy.
We split the sphere into two hemispheres. Without loss of generality, assume $T=p_{1}=(0,0,1)$, and that the hemisphere containing $p_{1}$ has $n_{1}\geq\frac{n}{2}$ robots. We move $p_{0}$ toward $p_{1}$ in $1$ time unit. Then, using the mapping $\mathcal{M}$, we map the robots on the boundary of the hemisphere containing $p_{1}$ onto an $l_{2}$-disk with radius $\frac{\pi}{2}$. We have $n_{1}$ asleep robots and one awake robot at the origin of the disk. Using the Arc-Strategy from Lemma 3.1.1, we can wake all the robots in the disk in $\frac{\pi}{2}\times 3.9851+2r_{1}$ time units, where $r_{1}$ is the distance of the nearest point to the origin. Since both $p_{0}$ and $p_{1}$ are at the origin, $r_{1}=0$, and all robots on the disk can be woken in $\frac{\pi}{2}\times 3.9851$ time units.
Thus, by Lemma 5.1 and the Arc-Strategy, after $1$ time unit (for moving $p_{0}$ toward $p_{1}$) plus $\frac{\pi}{2}\times 3.9851$ time units, all robots in the upper hemisphere are awake, giving us $n_{1}+1$ awake robots. Since $n_{1}\geq\frac{n}{2}$, each awake robot on the boundary can wake a robot in the lower hemisphere within $2$ time units. Therefore, all robots can be woken in $3+\frac{\pi}{2}\times 3.9851\simeq 9.26$ time units.
@@ 5.3. Our strategy for surface-FTP
On the other hand the Euclidean distance between their mapping $\mathcal{M}(p_{1})=p_{1}^{\prime}$ and $\mathcal{M}(p_{2})=p_{2}^{\prime}$ is:
@@
We claim that the geodesic distance between points $p_{1}$ and $p_{2}$ is less than or equal to the Euclidean distance between the corresponding points $p^{\prime}_{1}$ and $p^{\prime}_{2}$. To support this, we conducted a computational analysis, calculating distances for all combinations of $\delta_{1}=\frac{\pi}{2}\epsilon\times i$, $\delta_{2}=\frac{\pi}{2}\epsilon\times j$, and $\theta_{1}-\theta_{2}=\pi\epsilon\times k$, with $i,j,k$ ranging from 00 to $\frac{1}{\epsilon}$. The code, available online  [^2] Geodesic to Euclidean Distance Ratio Calculator Code, was run with $\epsilon=0.001$, and in all cases, the geodesic distance between $p_{1}$ and $p_{2}$ was less than or equal to $\|p^{\prime}_{1}-p^{\prime}_{2}\|$ [^3] A detailed computation of the derivatives of the distance functions can be found in our code..
Thus, the wake-up time for robots on the hemisphere is less than or equal to the wake-up time for robots on the disk. This means that a wake-up tree for robots on the disk also works for robots on the hemisphere, with a makespan that is no greater.
@@
  • We move $p_{0}$ to the nearest point, $p_{1}$, in $\rho_{1}$ time units, where $\rho_{1}$ is the geodesic distance between $p_{0}$ and $p_{1}$. Now, with two awake robots, $p_{1}$ is responsible for waking robots in the upper hemisphere, while $p_{0}$ handles the robots in the lower hemisphere.
  • Here’s the strategy for waking up the robots in the upper hemisphere: Using the mapping $\mathcal{M}$, we map the points on the upper hemisphere onto a disk with radius $\frac{\pi}{2}$, placing the awake robot at $p_{1}=(\rho_{1},\theta_{1})$. We then move $p_{1}$ to the center of the disk. The remaining robots on the disk can be awakened in $\frac{\pi}{2}\times 5.064$ time units using Theorem 2. Therefore, all robots on the upper hemisphere are awakened in at most $2\rho_{1}+\frac{\pi}{2}\times 5.064$ time units. Since $\rho_{1}\leq\frac{\pi}{2}$, the makespan for the upper hemisphere is at most $11.1$ time units.
  • Here’s the strategy for waking up the robots in the lower hemisphere: First, $p_{0}$ moves to $p_{1}$ in $\rho_{1}$ time units, then to $p^{\prime}_{1}$ in $\pi-\rho_{1}-\rho^{\prime}_{1}$ time units, where $p^{\prime}_{1}$ is the nearest point to the lowest point in the hemisphere. Using the mapping $\mathcal{M}$, we map the lower hemisphere onto a disk with radius $\frac{\pi}{2}$. Now, with two awake robots, $p_{0}$ and $p^{\prime}_{1}$, within $\rho^{\prime}_{1}$ of the disk’s center, the remaining robots can be awakened in $\frac{\pi}{2}\times 5.064$ time units using Theorem 2. The total makespan for the lower hemisphere is $\pi-\rho^{\prime}_{1}+\frac{\pi}{2}\times 5.064\leq 11.1$
Thus, after at most $11.1$ time units, all the robots are awake, and the wake-up ratio for surface-FTP is at most $11.1$.

### v4

@@ 4. FTP in $(\mathbb{R}^{3},l_{1})$
The figure shows how a cross-polytope with diameter d can be partiotioned intro six smaller cross-polytopes and eight smaller pyramids, each shape has diameter equal to d/2.
[…]
@@
[…]
A cross-polytope with diameter d is shown in the image, the shape is covered by six smaller crasspolytopes, these crosspolytopes overlap but cover the entire shape. The diameter of each small cross-polytope is equal to 2d/3.
@@ 4.1. Our strategy for FTP in $(\mathbb{R}^{3},l_{1})$
Given a set of $n$ asleep robots, our wake-up strategy is as follows: If $n\leq 127$, the awake robot at the origin can activate all robots within $13r$ time units, as shown in Lemma 4. If $n>127$, by Lemma 4, we first wake up 7 robots in $3r$ time units. Then, we divide the cross-polytope into six smaller overlapping cross-polytopes, each with a diameter of $\frac{2d}{3}$, as described in Lemma 4. With eight robots awake, we assign six of them to one of the smaller cross-polytopes. Each robot reaches the center of its assigned cross-polytope in at most $\frac{2d}{3}$ time units. Then, each robot wakes up the remaining robots in its section, taking at most $f\left(\frac{2d}{3}\right)$ time units, where $f(d)$ is the time needed to wake all robots in a region of diameter $d$. If a robot belongs to multiple cross-polytopes, it can be assigned to any one of them. If a cross-polytope has fewer than $128$ asleep robots, they can be woken up in $\frac{13d}{4}$ time units as per Lemma 4. Thus, the recursive formula for computing $f(d)$ is:
@@
Which yields: $f(d)\leq\frac{13d}{6}\times\left(1+\frac{2}{3}+\frac{4}{9}+\ldots\right)=\frac{13d}{2}.$ For the unit $l_{1}$-ball where $d=2$, the upper bound for the total time is $13$ units.
@@ 5.1. Mapping
Consider the upper hemisphere of a unit $l_{2}$-ball with $T=(0,0,1)$. Any point $p=(x,y,z)$ on the surface of the hemisphere can be uniquely represented by the pair $(\delta,\theta)$, where $\delta=\arccos(z)$ is the geodesic distance from $T$ to $p$ (the shortest path along the hemisphere), and $\theta$ is the angle of $p$’s projection onto the $xy$-plane relative to the positive $x$-axis. This is illustrated in Figure 8.
@@
$x$$y$$z$$\theta$$\delta$$T$$p$$\delta$$p^{\prime}$$r=\frac{\pi}{2}$${\theta}$$x$$y$
@@
[…]
The method of mapping a hemisphere onto a circle. Each point has some arc connecting it to the top of the hemisphere, the length of this arc is their geodesic distance and would be equal to the distance of the mapped point to the center of the mapped circle. The angle of the projection of this point on the xy plane with the x axis is directly mapped to the angle of the mapped point with the x axis, preserving angular position.
@@
[…]
The cases of the positions of two points relative to each other, figure a shows the case where both points have the same angular positions, figure b shows the case where both points have the same elevation. Figure c elaborates the general case by using the combination of the two previous cases, we create a regular trapezoid by using additional points.
@@
Now, we show that $\|p_{1}-p_{2}\|_{2}\leq\|p^{\prime}_{1}-p^{\prime}_{2}\|_{2}$. Consequently we need to show,
@@ 5.2. Our strategy, when asleep robots are on the boundary on the unit $l_{2}$-ball
For an instance of FTP in $(\mathbb{R}^{3},l_{2})$, where the asleep robots are on the boundary and the initially active robot is at the center of the ball, we propose the following strategy. We split the sphere into two hemispheres. Without loss of generality, assume $T=p_{1}=(0,0,1)$, and that the hemisphere containing $p_{1}$ has $n_{1}\geq\frac{n}{2}$ robots. We move $p_{0}$ toward $p_{1}$ in $1$ time unit. Then, using the mapping $\mathcal{M}$, we map the robots on the boundary of the hemisphere containing $p_{1}$ onto an $l_{2}$-disk with radius $\frac{\pi}{2}$. We have $n_{1}$ asleep robots and one awake robot at the origin of the disk. Using the Arc-Strategy from Lemma 3.1.1, we can wake all the robots in the disk in $\frac{\pi}{2}\times 3.9651+2r_{1}+2r_{2}$ time units, where $r_{1}$ is the distance of the nearest point to the origin and $r_{2}$ depends on the path and we have $r_{2}\leq 1$
. Since both $p_{0}$ and $p_{1}$ are at the origin, $r_{1}=0$, and all robots on the disk can be woken in $\frac{\pi}{2}\times 5.9651$ time units.
Thus, by Lemma 5.1 and the Arc-Strategy, after $1$ time unit (for moving $p_{0}$ toward $p_{1}$) plus $\frac{\pi}{2}\times 5.9651$ time units, all robots in the upper hemisphere are awake, giving us $n_{1}+1$ awake robots. Since $n_{1}\geq\frac{n}{2}$, each awake robot on the boundary can wake a robot in the lower hemisphere within $2$ time units. Therefore, all robots can be woken in $3+\frac{\pi}{2}\times 5.9651\simeq 12.37$ time units.
@@ 5.3. Our strategy for surface-FTP
On the other hand, the Euclidean distance between their mapping $\mathcal{M}(p_{1})=p_{1}^{\prime}$ and $\mathcal{M}(p_{2})=p_{2}^{\prime}$ is:
@@
We claim that the geodesic distance between points $p_{1}$ and $p_{2}$ is less than or equal to the Euclidean distance between the corresponding points $p^{\prime}_{1}$ and $p^{\prime}_{2}$. To support this, we conducted a computational analysis, calculating distances for all combinations of $\delta_{1}=\frac{\pi}{2}\epsilon\times i$, $\delta_{2}=\frac{\pi}{2}\epsilon\times j$, and $\theta_{1}-\theta_{2}=\pi\epsilon\times k$, with $i,j,k$ ranging from 00 to $\frac{1}{\epsilon}$. The code, available online  [^2] Geodesic to Euclidean Distance Ratio Calculator Code, was run with $\epsilon=0.001$, and in all cases, the geodesic distance between $p_{1}$ and $p_{2}$ was less than or equal to $\|p^{\prime}_{1}-p^{\prime}_{2}\|_{2}$  [^3] A detailed computation of the derivatives of the distance functions can be found in our code. Thus, the wake-up time for robots on the hemisphere is less than or equal to the wake-up time for robots on the disk. This means that a wake-up tree for robots on the disk also works for robots on the hemisphere, with a makespan that is no greater.
@@
  • We move $p_{0}$ to the nearest point, $p_{1}$, in $\rho_{1}$ time units, where $\rho_{1}$ is the geodesic distance between $p_{0}$ and $p_{1}$. Now, with two awake robots, $p_{1}$ handles robots in the upper hemisphere, while $p_{0}$ handles the robots in the lower hemisphere.
  • The strategy for waking up the robots in the upper hemisphere: Using the mapping $\mathcal{M}$, we map the points on the upper hemisphere onto a disk with radius $\frac{\pi}{2}$, placing the awake robot at $p_{1}=(\rho_{1},\theta_{1})$. We then move $p_{1}$ to the center of the disk. The remaining robots on the disk can be awakened in $\frac{\pi}{2}\times 5.4162$ time units using Theorem 2. Therefore, all robots on the upper hemisphere are awakened in at most $2\rho_{1}+\frac{\pi}{2}\times 5.4162$ time units. Since $\rho_{1}\leq\frac{\pi}{2}$, the makespan for the upper hemisphere is at most $11.65$ time units.
  • The strategy for waking up the robots in the lower hemisphere: First, $p_{0}$ moves to $p_{1}$ in $\rho_{1}$ time units, then to $p^{\prime}_{1}$ in $\pi-\rho_{1}-\rho^{\prime}_{1}$ time units, where $p^{\prime}_{1}$ is the nearest point to the lowest point in the hemisphere. Using the mapping $\mathcal{M}$, we map the lower hemisphere onto a disk with radius $\frac{\pi}{2}$. Now, with two awake robots, $p_{0}$ and $p^{\prime}_{1}$, within $\rho^{\prime}_{1}$ of the disk’s center, the remaining robots can be awakened in $\frac{\pi}{2}\times 5.4162$ time units using Theorem 2. The total makespan for the lower hemisphere is $\pi-\rho^{\prime}_{1}+\frac{\pi}{2}\times 5.4162\leq 11.65$
Thus, the wake-up ratio for surface-FTP is at most $11.65$.